# Thermal Inertia Controls on Titan’s Surface Temperature and Planetary Boundary Layer Structure

### v1

@@ 6 Discussion and Conclusions
At diurnal timescales, surface temperature amplitudes are highly sensitive to thermal inertia. Low thermal inertia surfaces respond inefficiently to subsurface heat conduction, leading to larger temperature variations and stronger sensible heat fluxes. In contrast, high thermal inertia surfaces store and conduct heat more effectively, resulting in reduced temperature variation amplitude. Theoretical predictions developed in this work accurately reproduce the simulated diurnal temperature variations across different cases, confirming that subsurface heat conduction is the primary damping mechanism at short timescales.
@@
At seasonal timescales, surface temperature variations show much weaker sensitivity to thermal inertia. Only when thermal inertia is uniformly elevated (e.g., Case 3) does seasonal temperature amplitude decline significantly. This is because the product , which characterizes the impact of subsurface heat conduction, is small compared to the atmospheric damping coefficient except under globally high thermal inertia. Consequently, seasonal surface temperature variations are mainly regulated by direct surface–atmosphere energy exchange through sensible heat flux, rather than by subsurface thermal diffusion.
PBL behavior differs markedly between diurnal and seasonal timescales. At diurnal timescales, the PBL exhibits a regular cycle of growth and decay—rising from approximately 500 [m] in the morning to nearly 1,000 [m] in the afternoon—driven by daily variations in surface heating and cooling. At seasonal timescales, the PBL responds to broader shifts in insolation. As thermal inertia increases, the contrast between minimum and maximum surface temperature diminishes, leading the seasonal PBL to extend to a maximum depth of 2,000–3,000 [m], consistent with previous dry GCM simulations (Charnay & Lebonnois, 2012), dune spacing analysis (Lorenz et al., 2010), and the dry adiabatic lapse rate extending to 3.5 [km] observed by Voyager IRIS (Lindal et al., 1983). These results highlight that PBL depth is governed not only by transient energy inputs but also by the long-term thermal state of the surface and overlying atmosphere.
@@
Simulated winds on Titan exhibit minimal diurnal variability, which is probably linked with the small diurnal temperature swings ( [K]) compared to Earth. However, the GCM accurately captures seasonal wind structures, including the key zonal and meridional features observed by the Huygens probe. The simulations support the interpretation by Tokano (2007) that Titan’s wind profiles are shaped primarily by global-scale circulation patterns, not Ekman dynamics as on Earth. A notable seasonal transition in the meridional wind is observed in the simulations: during southern summer, the flow is southward near the surface and northward aloft (Figure 7), while during southern winter, this pattern is reversed, with northward flow near the surface and southward aloft (not shown). This transition is characteristic of Titan’s cross-equatorial Hadley circulation.
@@
At the Dragonfly landing site, simulations during northern midwinter indicate a shallower diurnal PBL (reaching 500 [m]) due to reduced insolation compared to the Huygens site’s midsummer conditions. Nonetheless, temperature structures remain similar, and large-scale wind patterns are preserved, reflecting the shared equatorial latitude and the dominance of global atmospheric dynamics. These results suggest that Dragonfly will likely experience steady boundary layer conditions and modest winds, comparable to those encountered by Huygens.
@@
While our findings offer valuable insights, a key limitation of this study is the exclusion of methane moisture processes, which are likely to play a significant role in Titan’s climate system. Observations from the Huygens probe indicated that the surface was damp at the time of landing (Zarnecki et al., 2005; Lorenz et al., 2006), and our dry simulations overestimate surface temperatures by approximately 2 [K]. Furthermore, in several simulation cases, seasonal maxima in surface temperature and PBL depth occur at high latitudes, contrary to observational evidence indicating that Titan’s warmest regions are consistently located within 30° of the equator, as shown by Voyager IRIS data (Flasar et al., 1981; Flasar, 1998) and 13 years of Cassini CIRS observations (Jennings et al., 2016, 2019). Previous studies have demonstrated that the inclusion of moist processes substantially reduces the pole-to-equator surface temperature contrast, bringing simulated climate patterns into closer agreement with observations (Lora et al., 2019).
@@
This discrepancy highlights the importance of including latent heat processes from methane evaporation and condensation. Prior studies have shown these to play a critical role in regulating Titan’s surface energy balance and atmospheric structure (Griffith et al., 2008; Mitchell et al., 2009; Lora & Ádámkovics, 2017). Incorporating methane-driven moist processes into the surface energy balance term in Eq. (1) will be essential for future modeling efforts. Moreover, Huygens measurements of nearly uniform methane mole fractions in the lowest 4–5 [km] (Niemann et al., 2005, 2010; Gautier et al., 2024) suggest vigorous vertical mixing, possibly linked to the same large-scale circulation that shapes meridional wind reversals. All of this further motivates the inclusion of humidity and moist effects in future Titan PBL simulations.
@@
In summary, this study establishes a physically grounded framework for understanding Titan’s surface temperature and boundary layer variability across diurnal and seasonal timescales. By integrating theoretical energy balance analysis with dry GCM simulations, we demonstrate the important roles of subsurface conduction and atmospheric damping in modulating surface conditions. While future work must incorporate methane-related processes to capture Titan’s full climate complexity, our results provide critical constraints for interpreting Huygens-era observations and offer quantitative projections for environmental conditions at the Dragonfly landing site, supporting mission planning and future data analysis.
@@ Appendix A Theoretical Framework for Surface Energy Balance and Temperature Variability
where represents the complex amplitude of the temperature variation. The partial derivative of with respect to time is:
Substituting this into (A1) gives:
Assuming is constant (homogeneous subsurface), the equation simplifies to:
Rearranging it, we find:
@@
At , this becomes:
Applying the boundary condition (A2):
Rearranging gives:
@@
where [Wm-2] denotes the amplitude of fluctuations in . This expression is notably independent of the specific value of the deep boundary condition , ensuring its general applicability.

### v2

@@ 6 Discussion and Conclusions
Surface temperature amplitudes are highly sensitive to thermal inertia at diurnal timescales. Low thermal inertia surfaces respond inefficiently to subsurface heat conduction, leading to larger temperature variations and stronger sensible heat fluxes. In contrast, high thermal inertia surfaces store and conduct heat more effectively, resulting in reduced temperature variation amplitude. Theoretical predictions developed in this work accurately reproduce the simulated diurnal temperature variations across different cases, confirming that subsurface heat conduction is the primary damping mechanism at short timescales. At seasonal timescales, surface temperature variations show much weaker sensitivity to thermal inertia. Only when thermal inertia is uniformly elevated (e.g., Case 3) does seasonal temperature amplitude decline significantly. This is because the product , which characterizes the impact of subsurface heat conduction, is small compared to the atmospheric damping coefficient except under globally high thermal inertia. Consequently, seasonal surface temperature variations are mainly regulated by direct surface–atmosphere energy exchange through sensible heat flux, rather than by subsurface conduction.
Thermal inertia controls on the PBL are clearly evident in the simulations. At diurnal timescales, the PBL undergoes a regular cycle of growth and decay—rising from a few hundred meters during the nighttime to over 1,000 meters in the afternoon—driven by daily variations in surface heating and cooling. Surfaces with lower thermal inertia exhibit larger diurnal fluctuations in PBL depth, accompanied by stronger surface temperature variability. Notably, despite stronger daytime sensible heat fluxes, lower thermal inertia surfaces have shallower daytime PBL depths. At night, the PBL is consistently shallower over lower thermal inertia surfaces due to rapid surface cooling. These features are all linked to nighttime temperature inversions, which inhibit the development of deeper boundary layers. At seasonal timescales, the PBL responds to broader variations in insolation. As with the diurnal cycle, PBL depth increases with thermal inertia, reaching maximum depths of over 2,000 m in the highest thermal inertia case. This is consistent with previous dry GCM simulations (Charnay & Lebonnois, 2012), dune spacing analyses (Lorenz et al., 2010), and the dry adiabatic lapse rate inferred from Voyager IRIS observations (Lindal et al., 1983). The influence of high thermal inertia on PBL structure is also evident in simulations incorporating a global thermal inertia map, where enhanced PBL depths are found over regions such as Xanadu.
@@
Simulated winds on Titan exhibit minimal diurnal variability, which is probably linked with the small diurnal temperature swings compared to Earth. The GCM accurately captures seasonal wind structures in the lowest few kilometers, including the key zonal and meridional features observed by the Huygens probe. The simulations support the interpretation by Tokano (2007) that Titan’s wind profiles are shaped primarily by global-scale circulation patterns, not Ekman dynamics as on Earth. A notable seasonal transition in the meridional wind is observed in the simulations: during southern summer, the flow is southward near the surface and northward aloft (Figure 7), while during southern winter, this pattern is reversed, with northward flow near the surface and southward aloft (not shown). This transition is characteristic of Titan’s cross-equatorial Hadley circulation.
@@
At the Dragonfly landing site, simulations during northern midwinter show a shallower diurnal PBL—ranging from about 200 to 600 meters—due to reduced insolation compared to the midsummer conditions at the Huygens site. Nonetheless, temperature structures remain similar, and large-scale wind patterns are preserved, reflecting the shared equatorial latitude and the dominance of global atmospheric dynamics. These results suggest that Dragonfly will likely experience steady boundary layer conditions and modest winds, comparable to those encountered by Huygens.
A previous study estimated the maximum winds (i.e., the worst-case scenario that Dragonfly might encounter) associated with dry convective vortices using the equation , obtaining a value of 2.7 ms-1 with K (Lorenz, 2021). This estimate was based on a simplified energy balance approach in which the atmospheric temperature, , is held constant, similar to that of Williams et al. (2012). In our simulations, which explicitly account for the full energy balance, applying Eq. (7) with ms-1 as in Lorenz (2021) yields (Wm-2K-1). Combined with a sensible heat flux ranging from 0.5 to 1.0 Wm-2 as shown in Figure 8, this results in to K and consequently to ms-1, which is in good agreement with Lorenz (2021).
@@
While our findings offer valuable insights, a key limitation of this study is the exclusion of methane moisture processes, which are likely to play a significant role in Titan’s climate system. Observations from the Huygens probe indicated that the surface was damp at the time of landing (Zarnecki et al., 2005; Lorenz et al., 2006), and our dry simulations overestimate surface temperatures by approximately 2 K. Furthermore, in several simulation cases, seasonal maxima in surface temperature and PBL depth occur at high latitudes, contrary to observational evidence indicating that Titan’s warmest regions are consistently located within 30° of the equator, as shown by Voyager IRIS data (Flasar et al., 1981; Flasar, 1998) and 13 years of Cassini CIRS observations (Jennings et al., 2016, 2019). Previous studies demonstrate that the inclusion of moist processes leads to improved latitudinal profiles of surface temperature and more realistic zonal wind structures, thereby bringing simulated climate patterns into closer agreement with observations (Lora et al., 2015; Mitchell & Lora, 2016; Turtle et al., 2018; Lora et al., 2019).
@@
This highlights the importance of including latent heat processes from methane evaporation and condensation. Prior studies have shown these to play a critical role in regulating Titan’s surface energy balance and atmospheric structure (Griffith et al., 2008; Mitchell et al., 2009; Lora & Ádámkovics, 2017). Incorporating methane-driven moist processes into the surface energy balance term in Eq. (1) will be essential for future modeling efforts. Moreover, Huygens measurements of nearly uniform methane mole fractions in the lowest 4–5 km (Niemann et al., 2005, 2010; Gautier et al., 2024) suggest vigorous vertical mixing, possibly linked to the same large-scale circulation that shapes meridional wind reversals. All of this further motivates the inclusion of humidity and moist effects in future Titan PBL simulations.
@@
In summary, this study establishes a physically grounded framework for understanding Titan’s surface temperature and boundary layer variability across diurnal and seasonal timescales. By integrating theoretical energy balance analysis with dry GCM simulations, we demonstrate the important roles of subsurface conduction and atmospheric damping in modulating surface conditions and PBL structures. While future work must incorporate methane-related processes to capture Titan’s full climate complexity, our results provide critical constraints for interpreting Huygens-era observations and offer quantitative projections for environmental conditions at the Dragonfly landing site, supporting mission planning and future data analysis.
@@ Appendix A Theoretical Framework for Surface Energy Balance and Temperature Variability
where represents the complex amplitude of the temperature variation. Substituting this into (A1) and assuming is constant (homogeneous subsurface) gives:
@@
At , and using the boundary condition (A2), we obtain
@@
where (Wm-2) denotes the amplitude of fluctuations in . This expression is notably independent of the specific value of the deep boundary condition , ensuring its general applicability.